OpenStudy (anonymous):

Find tan θ if sec θ = square root 26/5 and sin θ < 0.

OpenStudy (anonymous):

So \[\sec \theta =\sqrt{26}/5\]

OpenStudy (anonymous):

Is it -1/5?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

that means \[\cos(\theta)=\frac{5}{\sqrt{26}}\]

OpenStudy (misty1212):

oh yeah think you are correct

OpenStudy (misty1212):



OpenStudy (misty1212):

the 1 by pythagoras, and tangent is negative, so you are right

OpenStudy (anonymous):

Thank you!

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]